# Bias in Gini coefficient estimation for gamma mixture populations

### v1

@@ 1 Introduction
Mathematically, the Gini coefficient is typically defined in terms of the expected absolute differences between two independent, identically distributed (i.i.d.) random variables. A widely used estimator for this index is the sample Gini coefficient; see Deltas, (2003). However, this estimator tends to be biased downward, particularly in small samples and across various distributions, including uniform, log-normal, and exponential models. To address this issue, Deltas, (2003) introduced the upward-adjusted sample Gini coefficient, which was shown to be unbiased when the population followed an exponential distribution. Recenty, Baydil et al., (2025) extended the unbiasedness of the upward-adjusted sample Gini coefficient when the population follows a gamma distribution.
@@
In this paper, we examine the estimation of the Gini coefficient in gamma mixture populations, a more general class of distributions that accounts for heterogeneity in income data. While the sample Gini coefficient is known to be an unbiased estimator for the population Gini coefficient in single gamma distributions (Baydil et al.,, 2025), we show that this property does not hold for gamma mixture models, leading to a systematic bias in estimation. Then, we derive an explicit expression for the bias and propose an unbiased Gini estimator when the population is gamma mixture distribuited.
@@ 2 Preliminary results and some definitions
Let $X_{1}$ and $X_{2}$ be two independent copies of a positive, absolutely continuous random variable $X$ with finite integral and common cumulative distribution function $F$ and let $g$ be a positive real-valued integrable function of two (positive) variables. Then, the following identity holds:
@@
The Gini coefficient (Gini,, 1936) for the random variable $X$ is defined as
@@
where $X_{1}$ and $X_{2}$ are two independent copies of $X$.
@@
where $m\in\mathbb{N}$ denotes the number of components, $\pi_{i}$ denotes the mixing proportion such that $\sum_{i=1}^{m}\pi_{i}=1$, $\pi_{i}>0$, and $g_{i}$ is the density of the gamma distribution with rate parameter $\alpha_{i}>0$ and rate parameter $\lambda>0$, that is,
@@
$G={\Gamma(\alpha+{1\over 2})\over\sqrt{\pi}\alpha\Gamma(\alpha)}.$
@@ 3 The main result
The following theorem is valid only for positive random variables and its proof adapts similar technical steps as reference Baydil et al., (2025).
@@
Let $X_{1},X_{2},\ldots$ be independent copies of a positive, absolutely continuous random variable $X$ with finite integral and common cumulative distribution function $F$. The following holds:
@@
and $\mathscr{L}_{F}(p)=\int_{0}^{\infty}\exp(-pu){\rm d}F(u)$ is the Laplace transform correspondig to distribution $F$. In the above, we are assuming that the Lebesgue-Stieltjes integrals and improper integrals involved exist.
@@
Under the conditions of Theorem 3.1, with $X$ being a positive, absolutely continuous random variable with finite integral and common cumulative distribution function $F$, we have
@@ 4 The Gini coefficient estimator is biased for gamma mixtures
where the well-known indentity $\Gamma(x)=\lim_{y\to\infty}\gamma(x,y)$ has been used. Furthermore, since $X^{*}\sim\text{GM}(\bm{\theta}^{*})$ with $\bm{\theta}^{*}$ being as in Example 2.1, by using (4), we have
@@
$-\sum_{1\leqslant i_{1},\ldots,i_{n-1}\leqslant m}\pi_{i_{1}}\cdots\pi_{i_{n-1}}\,\sum_{j=1}^{m}{n\pi_{j}\alpha_{j}\over\alpha_{i_{1}}+\cdots+\alpha_{i_{n-1}}+\alpha_{j}},$
@@
Therefore, the bias of $\widehat{G}$ relative to $G$, denoted by $\text{Bias}(\widehat{G},G)$, can be written as
@@ 5 Illustrative simulation study
Figure 2 shows the average values of the standard and bias-corrected Gini coefficient estimates for different sample sizes $n$, holding $\alpha_{1}=0.5$ and $\alpha_{2}=2.0$ constant. From this figure, we observe that the standard Gini coefficient estimator ($\widehat{G}$, red line) consistently overestimates the true Gini coefficient. However, as the sample size increases, the bias diminishes slightly, demonstrating a convergence towards the true value, as expected. The bias-corrected estimator ($\widehat{G}_{\text{bc}}$, blue line), on the other hand, consistently produces more accurate estimates across all sample sizes.
@@ 6 Concluding remarks
In this paper, we have proposed the bias in the estimation of the Gini coefficient for gamma mixture populations. While the sample Gini coefficient is unbiased for single gamma distributions, we have demonstrated both theoretically and empirically that this property does not hold for gamma mixture models. We have derived a bias expression for the Gini coefficient for gamma mixture populations, which allowed us to propose a bias-corrected Gini estimator. An illustrative Monte Carlo simulation study has been carried out to evaluate the behavior of the bias-corrected Gini estimator. The results emphasized the that the standard Gini coefficient estimator exhibits a consistent upward bias when applied to gamma mixture populations, and the proposed bias correction effectively mitigates this issue, providing a more reliable estimator for gamma mixture distributions. As part of future research, it will be of interest to extend the study to multivariate Gini coefficients. Futhermore, alternative correction methods or the analysis to other mixture models can be developed; see Pérez et al., (1986). Work on these problems is currently in progress and we hope to report these findings in future.

### v2

@@ 1 Introduction
Mathematically, the Gini coefficient is commonly defined in terms of the expected absolute differences between two independent, identically distributed (i.i.d.) random variables. A widely used estimator for this index is the sample Gini coefficient; see Deltas, (2003). However, this estimator tends to be biased downward, particularly in small samples and across various distributions, including uniform, log-normal, and exponential models. To address this issue, Deltas, (2003) introduced the upward-adjusted sample Gini coefficient, which was shown to be unbiased when the population followed an exponential distribution. Recently, Baydil et al., (2025) extended the unbiasedness of the upward-adjusted sample Gini coefficient when the population follows a gamma distribution.
@@
In this paper, we examine the estimation of the Gini coefficient in gamma mixture populations, a more general class of distributions that accounts for heterogeneity in income data. While the sample Gini coefficient is known to be an unbiased estimator for the population Gini coefficient in single gamma distributions (Baydil et al.,, 2025), we show that this property does not hold for gamma mixture models, leading to a systematic bias in estimation. Then, we derive an explicit expression for the bias and propose an unbiased Gini estimator when the population is gamma mixture distributed.
@@ 2 Preliminary results and some definitions
Let $X_{1}$ and $X_{2}$ be two independent copies of a (almost surely) positive random variable $X$ with finite integral and common cumulative distribution function $F$ and let $g$ be a positive real-valued integrable function of two (positive) variables. Then, the following identity holds:
@@
The Gini coefficient is defined using the Lorenz curve, which plots cumulative income share against population percentage. The line of perfect equality is a 45-degree line. The Gini coefficient is calculated as the ratio of the area between the Lorenz curve and the line of equality to the total area under the line. To derive explicit expressions, we use the standard definition based on the mean difference.
@@
The Gini coefficient (Gini,, 1936) of a random variable $X$ with finite mean $\mathbb{E}(X)$ is defined as
@@
where $X_{1}$ and $X_{2}$ are independent copies of $X$.
@@
where $m\in\mathbb{N}$ denotes the number of components, $\pi_{i}$ denotes the mixing proportion such that $\sum_{i=1}^{m}\pi_{i}=1$, $\pi_{i}>0$, and $g_{i}$ is the density of the gamma distribution with shape parameter $\alpha_{i}>0$ and rate parameter $\lambda>0$, that is,
@@
$G={\Gamma(\alpha+{1\over 2})\over\sqrt{\pi}\alpha\Gamma(\alpha)},$
@@ 3 The main result
The following theorem is valid only for (almost surely) positive random variables and its proof adapts similar technical steps as reference Baydil et al., (2025).
@@
Let $X_{1},X_{2},\ldots$ be independent copies of a (almost surely) positive random variable $X$ with finite integral and common cumulative distribution function $F$. The following holds:
@@
and $\mathscr{L}_{F}(p)=\int_{0}^{\infty}\exp(-pu){\rm d}F(u)$ is the Laplace transform corresponding to distribution $F$. In the above, we are assuming that the Lebesgue-Stieltjes integrals and improper integrals involved exist.
@@
Under the conditions of Theorem 3.1, with $X$ being a (almost surely) positive, absolutely continuous random variable with finite integral and common cumulative distribution function $F$, we have
@@ 4 The Gini coefficient estimator is biased for gamma mixtures
where the well-known identity $\Gamma(x)=\lim_{y\to\infty}\gamma(x,y)$ has been used. Furthermore, since $X^{*}\sim\text{GM}(\bm{\theta}^{*})$ with $\bm{\theta}^{*}$ being as in Example 2.1, by using (4), we have
@@
$-\sum_{1\leqslant i_{1},\ldots,i_{n-1}\leqslant m}\pi_{i_{1}}\cdots\pi_{i_{n-1}}\,\sum_{j=1}^{m}{n\pi_{j}\alpha_{j}\over\alpha_{i_{1}}+\cdots+\alpha_{i_{n-1}}+\alpha_{j}},$ (26)
@@
Therefore, by (26) and (9), the bias of $\widehat{G}$ relative to $G$, denoted by $\text{Bias}(\widehat{G},G)$, can be written as
@@ 5 Illustrative simulation study
Figure 2 shows the average values of the standard and bias-corrected Gini coefficient estimates for different sample sizes $n$, holding $\alpha_{1}=0.5$ and $\alpha_{2}=2.0$ constant. From this figure, we observe that the standard Gini coefficient estimator ($\widehat{G}$, red line) tends to overestimate the Gini coefficient (positive bias). However, as the sample size increases, the bias diminishes slightly, demonstrating a convergence towards the true value, as expected. The bias-corrected estimator ($\widehat{G}_{\text{bc}}$, blue line), on the other hand, consistently produces lower estimates across all sample sizes.
@@ 6 Concluding remarks
In this paper, we have proposed the bias in the estimation of the Gini coefficient for gamma mixture populations. While the sample Gini coefficient is unbiased for single gamma distributions, we have demonstrated both theoretically and empirically that this property does not hold for gamma mixture models. We have derived a bias expression for the Gini coefficient for gamma mixture populations, which allowed us to propose a bias-corrected Gini estimator. An illustrative Monte Carlo simulation study has been carried out to evaluate the behavior of the bias-corrected Gini estimator. The results emphasized the that the standard Gini coefficient estimator exhibits a consistent upward bias when applied to gamma mixture populations, and the proposed bias correction effectively mitigates this issue, providing a more reliable estimator for gamma mixture distributions. As part of future research, it will be of interest to extend the study to multivariate Gini coefficients. Furthermore, alternative correction methods or the analysis to other mixture models can be developed; see Pérez et al., (1986). Work on these problems is currently in progress and we hope to report these findings in future.